Kevin has money in two savings accounts. One rate is 6% and the other is 12%. If he has $600 more in the 12% account and the total interest is $279, how much is invested in each savings account?
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52943)

.
Kevin has money in two savings accounts. One rate is 6% and the other is 12%.
If he has $600 more in the 12% account and the total interest is $279,
how much is invested in each savings account?
~~~~~~~~~~~~~~

x = the amount saved at 12%;  

y = the amount saved at 6%.


From the condition, you have these two equations

    x     -     y = 600   dollars     (1)

    0.12x + 0.06y = 279   dollars     (2)


To solve this system, express x = y + 600 from equation (1) and substitute it into equation (2).  You will get

    0.12(y+600) - 0.06y = 279

    0.12y + 0.12*600 + 0.06y = 279

    0.12y + 0.06y            = 279 - 0.12*600

         0.18y               = 207

             y               = 207/0.18 = 1150.


Then from equation (1),

    x = 600 + 1150 = 1750.


ANSWER.  $1750 invested at 12%  and  $1150 invested at 6%.


CHECK.  I will check equation (2).   Its left side is  0.12*1750 + 0.06*1150 = 279  dollars.   ! Correct !

Solved.



Answer by greenestamps(13216)



There are numerous ways to set up this problem for solving using formal algebra; the other tutor has shown one.

If a formal algebraic solution is not required, you can get some good mental exercise by solving the problem informally, maybe something like this:

The "extra" $600 invested at 12% yields $72 interest.
The remaining amount, invested equally at 6% and 12%, gives an average rate of 9%; and it yields the remaining $207 interest.
$207 interest at 9% means the amount invested was $207/.09 = $2300.
That $2300 is equally divided between the two investments, so the amount invested in each was $1150.
Now remember the additional $600 invested at 12% interest to find that....

ANSWER: $1150 at 6%; $1150+$600 = $1750 at 12%

CHECK: .06(1150)+.12(1750) = 69+210 = 279